I need help learning how to solve percentage problms for the ASVAB i'm going to be taking. I don't how to begin solving a % problem. Can someone help me.
Answer by checkley79(3341)

What percentage is 7 of 28?
7/28=.25=25% ans.
------------------------
55 is what percentage of 22?
55/22=2.5=250% ans.
-------------------------
45% of 1440?
1440/45=32 ans.
-------------------------
.6% of 74?
.6*74=44.4 ans.
------------------------
Hope these examples helps.